If the length of a side of a square is increased by 3 centimeters, the perimeter of the new square is 40 centimeters more than twice the length of the side of the original square. Find the length of the side of the original square.
Answer by CharlesG2(834)

If the length of a side of a square is increased by 3 centimeters, the perimeter of the new square is 40 centimeters more than twice the length of the side of the original square. Find the length of the side of the original square.
squares have 4 equal sides by definition
the area A of a square equals side s squared or A = s^2
the perimeter P of a square equals 4 times side s or P = 4 * s
P = 2 * s + 40 = 4 * (s + 3)
2*s + 40 = 4*s + 12
2*s - 4*s = 12 - 40
-2 * s = - 28
s = 14 cm
check:
2*14 + 40 = 4 * (14 + 3)
28 + 40 = 4 * 17
68 = 68